OpenStudy (anonymous):

What is the 20th term of the sequence that begins -3,6,-12,24...? A.-524288 B.524,288 C.1,572,864 D. 3,145,728

OpenStudy (welshfella):

can you see a pattern here Hint: each term is obtained from the previous by multiplying by a sertain number - can you see which number?

OpenStudy (anonymous):

3?

OpenStudy (welshfella):

no 6 / -3 = ?

OpenStudy (anonymous):

-2??

OpenStudy (welshfella):

-12/6 = ?

OpenStudy (welshfella):

yes -2

OpenStudy (anonymous):

ah gotchaa! okk

OpenStudy (welshfella):

this is a geometric sequence and the formula for the nth term is a1*r^(n-1) where a1 = first term and r = common ratio

OpenStudy (welshfella):

in this case a1 = -3 and r = -2 so 20th term = -3 * (-2)^(20-1) = -3 * (-2)^19

OpenStudy (welshfella):

can you work that out - use calculator of course

OpenStudy (anonymous):

i got 1572864?? :S i feel like thts wrong

OpenStudy (welshfella):

no thats correct

OpenStudy (anonymous):

ahhh it is indeed ^.^ sorry. thank you!!

OpenStudy (welshfella):

yw